OpenStudy (anonymous):

Particular antiderivatives f(x) = x^5 - 2x^-2 + 1; F(1) = 0

OpenStudy (anonymous):

\[x^{5}-2x^{-2}+1\]

OpenStudy (anonymous):

i know the antiderivative is \[\frac{ x^6 }{ 6 } -\frac{ 2 }{ x }+x+C\]

OpenStudy (anonymous):

but when i try and find C i come out with -5/6 when the answer should be -19/6

OpenStudy (anonymous):

yeah that was the problem \(-\frac{2}{x}\) should be \(\frac{2}{x}\) with no minus sign that is why it didn't work

OpenStudy (anonymous):

ohhhhh i see the error now... thank you